mainmenu

DFR Telecoms Diary

2017

January 2017

1 January :

Welcome to 2017.

The jobs we know about for 2017 include :
1) Being organised about maintenance
2) Training
3) Installing Parkend North exchange and concentrator.
4) Moving DP3S into a new and more accessible cabinet


4 January : Charles and I went to the Junction to change the 50 volt batteries in the signal box. The new set appear to be good. Rick and Martin went to check the signal post phone at SP2. The PG was caused by the heavy steel handset cord preventing the handset from remaining on the cradle. Perhaps the cold was to blame, however we will consider changing the phone as the problem is likely tro persist. Ray and Peter went round checking clocks which somehow had got out of sync. Ray also went off to Parkend to test the main 50 volt battery by taking off the mains for a period. All OK. However he found junction 2 outgoing to be dis out, making level 6 and 0 calls unavailable. Rick and Martin went to Whitecroft to break the circuit, the fault is again in the Norchard to Whitecroft section. We will look for it next week. I became entry 1/2017 for an accident when I crushed my thumb between the hacksaw frame and the copper pipe I was cutting. It bled! However the resultant clean copper will fetch a nice price for our telecoms fund.

The new DP3S cabinet

7 January : Paul and I spent the day fixing the cabinet for DP3S to the wall by Betty's hut. We had to make a couple of trips to Light Fantastic to get backboard and floor wood and the wall fixings. We also had lunch at Kaplan's with Jean.

11 January : Rick and Martin spent their day looking for the cable fault on the Parkend junction. They narrowed it down to the section south of Whitecroft. They had to move to a spare pair to clear the problem. This leaves us with a suspect cable length. Peter and Charles spent their day disconnecting the C&W GUV and shifting the circuit to the workshop DP ready for a new sub DP to be provided when we can find someone prepared to climb a ladder. The GUV moves to the long siding beside the workshop on Saturday. Once in place it is proposed to erect a building over two of the sidings for stock protection. I tried to find out what the proposals are for the shop rearrangements. They look as though it will be physically difficult to provide the data sockets they will need. I will chase Jason over this and the associated electrical work. I managed to get the 20 pair cable to the new DP in the exchange fanned out on the MDF block but then got sidetracked onto a fault leaving three linefinders off normal with blown fuses. I then got sidetracked off this problem when I leaned on the signal post relay set to find it red hot. The uniselector inside was permanently operated. This came down to the residual on the controlling relay having been worn away so that the relay stuck on. A new armature with adjustable residual was fitted.

14 January : Rick, Peter and I were in. We started the day with discussions with Jason, Lloyd the new electrician, and James about the requirements for the reorganised shop. The counter has already arrived in the form of flatpack shelving units. They do not lend themselves to mains and data cabling. It seems that we will need to cut out holes between the various compartments through which to run cables and leads. Whatever we do, I'm sure there are some clever clogs out there who will explain what we did wrong! More discussion is needed when we can get to Lloyd, as he will have to take the lead in installing the cables and sockets. Rick spent time chasing a complex fault on the A2 linefinders. In the end no fault was found but it was all working well when we left. A 3 minute guarantee job. Peter worked on moving the BT analogue cordless phone from the counter to the shop office. I terminated Peter's 20 pair cable on the MDF.

18 January : Peter got his job in the museum quite a way forward with help from Charles and Ray. Rick and Martin went to see if they could locate the fault in the cable at Whitecroft with the echo tester. It's about 70 m from the Whitecroft cabinet which puts it in the area of the two epoxy filled joints on the fence line. I think that I may have to find a replacement length to cover the area of the fault, or better still, to replace the total section length. I helped just a little in the museum after finding out about the new purchase order system. After lunch we had a bit of a tidy round in the office.

21 January : It was the day of the Telecoms Group Lunch at our house. Jean and I welcomed Rick and Janet, Martin and Cynthia, Peter, Charles and Alison, Ray and June. Jean put on a good show of food.

25 January : Rick and Ray were both away. Martin made a start on sorting and thinning our small stores. He has been asked to be ruthless in throwing away items we have little use for. Peter and Charles carried on in the museum. I could do little as there was no room to work on the exchange frames, so I left after lunch expecting to do housework at home (all done when I got there).

Constructing a temporary coal pile

28 January : The railway was closed to the public today as Scott was digging out a large trench for a temporary coal pit. As the morning progressed we found the cable to the DP at the caboosh was dis. In the end we decided not to try to find the damage but to abandon the phones in our caboosh and Jason's hut nearby. The plans for the shop layout have changed but I could not find anyone with an up to date copy, so we will find out hopefully when we go in on Wednesday. The shop was in darkness with the electrician shifting the many light switches to the shop office. The shop was also completely empty with the stock now residing in the museum. Doors for the external corridor were being constructed and fitted. In the midst of the mayhem, Peter and Paul managed to get the Asterisk junctions rerouted via the new museum DP. We gave up at lunchtime and went to meet Jean for lunch at Kaplan's.

February 2017

1 February : The railway was still closed to the public today, apart from the cafe. Just as well as the site is mostly wet mud. The shop is a bare area with the hall doors now fitted and the internal wall removed. However we found that the shop layout has changed yet again and that no one on site knows what the current plan is. How very DFR! On our way out we also found that wireless CCTV is being considered for the shop. Yet more eggs in the BT router basket. We await Saturday's updates. We did get to recover the phones off DP12, reterminate the DP cable onto the block on the outside of the building and recover the cables between DP5A and the blocks in the Severn and Wye room. At least we have tidied something out of the cable cutting. After lunch we were introduced to Justin Eager, ex BT and wondering about volunteering with us. We showed him round a bit and had quite a chat. He may call in on Saturday morning to see what we do.

Peter was the last to leave the site and so he was the one called to a site meeting in the shop to discuss the provision of data sockets. In the end it seems that sockets are required on the outside wall of the shop office for use by the booking clerks and on the outside wall of the shop for use by the shop manager. Of course that is where it is most difficult to provide our cables and extremely difficult to provide electrical services, so a trunking route has been agreed with the planners. Saturday may be interesting.

4 February : Peter and I attended today, mostly to see what was the latest in the shop. Rick and his dog called in for tea. Plasterboard and plastering was going on and access to the exchange was difficult. The shop layout seems to be hardening, though we could not find out what the counter construction would be like nor exactly where they would be positioned. However surface trunking and a hole through the shop office wall seems to be agreed. Due to the uncertainty of the future we have decided to fit eight data sockets on each of the two counters so that whatever they get used for, there will be enough to cope. We still need to discuss this with the electrician who was not in today. We had a very slow lunch with Jean in a very busy Kaplan's. We managed to change out a faulty data insert in the cafe and repair a dis that we obviously put onto the adjacent sockets. Such a pity that the GPO had not designed the sockets, we would not be getting these problems. Peter and I also recovered the old cables to the asterisk and old museum DPs now that everything is working via the new DP.

8 February : We started the day with a tour of the shop. It is currently being painted. Part of the carpet is down but decisions about the counters are still awaited. We could hardly get to the exchange so Rick, Martin and Charles went mob handed to measure the length of the faulty cable section at Whitecroft. 361m point to point, say a replacement length of 400m of ten pair 0.9mm or larger if possible is required. Ray and I took the exchange meter readings, still lots of traffic to the Parkend and Lydney Junction exchanges. We had lunch together in the cafe and more chat and decided to go home early.

11 February : Peter, Rick and I were in for the morning. Currently we now expect that only a temporary counter will be installed for some time. Our electrician, Lloyd, and I agreed that we would not be able to do a permanent job until the final counters are installed. It may be necessary for James to work with temporary mains and data cables for some time. It may be quite a long time before we can get to finish the job. The window had been boarded up in the shop and we helped James remove the offending boarding. Rick pushed the old data cabling that was hanging out of the ceiling, back into the loft for now. We also measured and cut the backboard for the new DP3S and on the way home I bought the necessary paint. We picked up Jean and Janet before having lunch at Kaplan's.

15 February : Rick and Charles changed the phone at Middle Forge which had a dial that had seized after six months use. It was badly corroded but worked when WD40 was used on it. Peter worked on his 3 volt / 12 volt power supply for the museum D phones. Ray and I refitted the shop clock although it needed a good scrub to get the paint off of it. The shop has James refitting shelving etc but no decisions have been made about the counter yet. I took a box of copper to Bendall's and got a very good price for the scrap. At least it declutters the office a bit.

18 February : Paul and I were got at as soon as we arrived at the railway. The Parkend token was not working and the railway had changed to using a pilotman. No morning cup of tea for us, it was straight into checking the circuit. The phone at the Parkend ground frame was also reported out of order so we thought that the cable to the frame had been damaged. We went to Parkend to find that the phone was just sticky handset contacts. Perhaps we should go back and check them again. We checked the token and proved that we could send a battery to Norchard but could not receive one. We went to DP1 and proved the same condition there, conclusion, dis somewhere down the line or at the Norchard token machine. Returning to Norchard we proved the same condition out of DP3S, ie the dis was probably between Norchard site and the Parkend site. We went to Whitecroft to break down the circuit halfway to prove the fault towards Norchard. We tested the dis pair to find a problem 70 metres towards Norchard, ie the same as the previous fault. We changed out the pair by sacrificing the 597 line at the gate cabinet at Whitecroft. We returned to Norchard to make the same change at DP3S. The circuit then worked. We then went to meet Jean for lunch at Kaplan's having not had time for a cuppa all morning. After lunch we wondered what to do with the rest of a spoilt day and decided to check the Lydney Junction public address which had been reported as faulty. Despite a signalman who would not let us onto his newly polished floor we managed to look over the equipment and get the signalman to make a test announcement. It worked. Right when tested (RWT). I still do not want to be a maintenance man. Roger has been chased about getting hold of a length of suitable cable as we now have two pairs out of ten faulty. I wonder what Wednesday will bring to us. By the way, the shop is looking rather like it always has with the old counters back in their original positions. At present there is no decision on the new counters.

22 February : I had a training session with Martin and Charles about 3000 type relays. Rick was repairing the new DP3S cabinet. Peter and Ray were clearing up the last of the work via the museum DP and getting the records together. We were joined at lunch by Ian and his birthday cake.

25 February : Peter recovered unwanted jumpers and Paul connected the clock 70A to the series clock circuit. It didn't work, it needs much more current and will have to be fed from a separate feed. I checked our spare drum of 14 pair/0.9mm cable to find it in good condition and 425m in length. It looks as though we may have to use it at Whitecroft. Jason said he could move the drum to site and lay out the length, possibly in April. We had lunch at Kaplan's.

27 February : It was the day for the group to have an outing to an Indian restaurant in Gloucester. We had health problems affecting some of the group so in the end it was Peter (the Organiser), Martin and Cynthia, Rick and Janet, Charles and Alison who went for an excellent evening out.

March 2017

1 March : A pretty quiet day. Rick got on with selector maintenance. Peter checked three signal post phones we had inherited. They all ended up in the skip. Otherwise it was mostly a tidy around day.

3 March : Another group outing day, this time to the Forest of Dean College lunch. Probably the best meal there so far. We look forward to next month's lunch.

8 March : Ray and Rick changed the phone at Parkend ground frame. The existing phone was faulty due to damp and corrosion. The ground frame hut is very damp with no heating so we can expect more problems in the future. A junction from Lydney Signal Box was not working. I went to the Junction with Martin and Charles to start faulting. In the end it was proved into the cable network. As I had to leave early, the fault was handed over to Rick. After lunch Ray, Charles (not the singer) and Rick started to look at the fault on Lydney Junction Sig Box to Norchard junction 1. It was definitely dis one leg not far from Norchard. After a slight problem with the records, a visit to DP 3E revealed an HR conn on E side pr 8. A visit to the junction signal box found that the fault still existed. Downstairs we had problems proving which junction was faulty as I/c and O/g labels seem transposed. Back to Norchard and further testing indicated that Jcn 2 was now 1 leg dis. Another visit to DP 3E and we found a broken wire on D side 9 (probably from our previous visit). We will probably have more problems at this DP as it has a 20 pr box conn with IDC connexions with wires that are too thick. Test calls were passed and both junctions are working for now. Peter and Martin worked on DP13 and also fitted new trunking and 3 new cables from the 10/50 in the museum.

11 March : Just Paul and I in at the railway today. We had a quiet cup of tea in the cafe and considered the day. James though, just in passing, asked if we could provide a BT phone and an internal phone on the third desk now in the office for Andy Blake, our new General Manager. No one was in the office so moving items around, pulling off trunking covers and undoing sockets did not worry anyone. We got the BT phone working quite quickly on the end of a data link back to the shop cabinet. The internal phone was more awkward. Wired sockets existed but after getting the oscillator and amplifier working, we concluded that the cable had been cut off in the trunking. It had, and we now need to cable afresh back to DP14 in what was the booking office but is now the file room. By then our half day had ended. On the way out we had a thought. The clock cable above the desk probably has spare pairs. If so we have an easy job to get the phone to the desk. Of course, as soon as we have done that, "they" will take the wall down! We collected Jean for lunch at Kaplan's.

15 March : Rick and Martin worked on maintenance in the exchange. Charles continued museum work on the bells to be added to the token demonstrator. Ray surveyed for the upgrade work necessary on the PA system. We have two non working speakers and one now missing after the shop rearrangements. I met our new General Manager, Andy Blake, and had a chat with him, particularly over our worries concerning the data network. I spent most of the day cleaning up scrap wiring ready for sale at the best price at Bendall's. We all spent a few moments in the sun contemplating the inside of DP3E. It's becoming a fault liability as it was originally provided with some IDC connectors which are now looking pretty grotty. A quiet day, with no surprises for once.

18 March : Peter and I spent the morning installing the internal phone for the GM. A simple job made tedious by an incorrect record, however the tone tracer quickly told us what we needed to know. After lunch Peter and I met Steve Harris at the new Parkend signal box to discuss our cabling requirements. Paul spent the day with Sam Bolingbroke going over our data network. Sam seems likely to take the lead for future data work but will still require our help in cabling and other physical work etc.

19 March : We have received a very large parcel from Cornwall containing a tool box with its strowger tool kit. Lovely, it will be ideal to go to our new Parkend North exchange when we install it later this year. Thanks very much for the gift, Dave.

22 March : Rick got on with making the catch arrangements for the DP3S Cabinet door. Drilling, bending, cutting of substantial steel was involved. Peter and Martin continued to take over the exchange while moving museum circuits to the new DP. Ray occupied himself with getting down to detail on the planning for the PA speaker cabling upgrade we need to accomplish. I took the pile of scrap copper pipe we have been given and started to cut off the unwanted brass fittings. I also tidied up the pile of cable ends we have so that they ended up in our dustbin. I had intended to take the copper and cable to the scrap yard but the rain was so heavy I simply went home instead. Of course the sun came out once I had removed my wet coat and boots!

24 March : I took the scrap cable and copper that we had collected (since the 15th of February) to Bendall's and sold it for 85. Well worth the effort. Perhaps we can manage the PA speaker upgrade from the fund rather than charge the railway.

25 March : Peter, Paul and I caught the train to Parkend inorder to leave the concentrator head and phone in the signal box for S and T to fit. Whilst travelling there we were told that signal post 2 is still giving trouble and that a key was stuck in the lock of the platform phone at Parkend. We could not release the key and so reported the problem to the GM for attention by a locksmith. This resulted in us being asked to provide the GM with 00 as his phone number. Quickly jumpered up. However we now need to reissue the directory, so I have asked James to consider what internal phones will need to be provided with the new counter arrangement so that we do not have to reissue the directory again. Paul found that the signal post relay set in the exchange was red hot. The uniselector was permanently energised. So that's why the charge current was up and the exchange volts down. Peter and Paul went through the relay set checking residuals and armature travel etc so that the uniselector is now stepping very smartly indeed. They also drilled a hole in the side of the 10+50 to accept new cables from the new museum DP. At that point we decided to go to tea and finish the day. At one point we asked Paul if he knew where the red PA cable was. In the caboosh was the answer. Paul went to have a look and came back grinning. The cable was there all the time still on the cable dispenser which was in the middle of the floor. We must have moved it back and forth as we searched the caboosh. The cable did not come from Screwfix as we thought but was purchased from Canford Audio. Perhaps we have lost it.

29 March : After lunch I called briefly at the railway. Charles was busy with his bells for the museum token machines and Rick, Peter and Martin were continuing on with work on the museum exhibits. When I got home James rang to say that the SPT at the abutment was hanging off the wall. A job for Saturday.

April 2017

1 April : The railway was full of visitors as locals were given the opportunity for a free trip. Both the steam and diesel trains were packed. Paul, Peter and I trouped off up the line to the abutment where we found the SPT hanging rather loosely by two screws. We had to return to base for the necessary tools etc. Paul and I returned to refix the phone to the wall. It's now very firmly affixed. It tested fine to the signal box. We gave up on the cafe and went to meet Jean for lunch at Kaplan's. After lunch it rained and put us off the idea of checking the phone at the Parkend ground frame. It is supposed to be out of order yet again. Either no dial tone or a dial going at the wrong speed are suggestions from the Operations staff. Peter got back to consider the best way to further integrate the museum 10+50 into the network and we feel that getting extensions 1-10 working well will be the first aim. Then circuits can be transferred to them and the battery supply to the rest of the doll's eyes cut off so that they cease to operate incorrectly. Paul is thinking about "dial a disc". For a first trial he has connected the Thomas music to 416.

2 April : I got a call from James saying "no communication with Parkend". I checked that we could make outgoing calls to Parkend using my asterisk phone. We drove up to Parkend to find outgoing 1 seeming to be "dis". I busied the junction out and all was well on junction 2. Parkend is now back in touch, but no doubt the cable at Whitecroft is continuing to deteriorate.

5 April : Rick and I chased the fault on the Parkend junction. Still in the same cable length. Unfortunately all the spares are now faulty so that we could not rectify the problem. I have escalated the problem to both the General and Operational Managers and my Director so that they are aware that we could lose the token circuit at any time. It is now necessary to get Jason, our Development Manager, in the loop and hope he can find suitable staff to install the replacement cable quite quickly. Charles and Ray changed a dial in the signal box. Ray checked the stores necessary for the proposed PA connectors. Peter and Charles continued to work on the museum exhibits.

7 April : Rick and Janet, Martin and Cynthia, Peter, Ray and June, John and Jean attended a lunch at the Forest of Dean Catering College. Very good and excellent value.

I met Jason yesterday on site and we agreed that as the condition of the site stands and with the development proposals that exist, there is not a suitable route for a buried cable on either side of the line. Consequently Jason will seek approval to hire in a large machine to clear the spoil heaps on the site and clear a route close to the west hedge line. He will then trench and lay the cable direct in ground. This is very satisfactory from an S and T point of view. The cable will be safe and its position will be recorded. He hopes to proceed in the next few weeks.

The Longmoor Military Railway 060 Saddle Tank Rennes having a run out

12 April : Rick got back to fitting the door to the new DP3S cabinet. Ray and Charles dealt with phone faults at Norchard signal box and Whitecroft. Martin is back recording our small stores. We had a go at clearing junk out of the exchange at Norchard and I eventually bought a dustpan and brush to get the foor clean. The place looks a treat now. I made a small start on the caboosh, at least I can find the floor at the moment.

19 April : I met Steve Bradbury and two colleagues from the SVR and gave them a tour of the telecoms exchanges using the DMU to get about. Rick and Martin went to change the SPT at Sig Post 2 but eventually decided that it was the pair that was "dis". The fault locator gave a distance that did not seem to lead to a problem and in the end, time ran out and the fault will have to wait until next week. Ray and Charles are looking at the PA upgrade and getting the exchange end connectors fitted. We did get good news from Jason. He has the cable drum ready to move to site and the large machine he needs is likely to be available tomorrow.

21 April : Ray was called in to find all the reverse pulse clocks stopped, loop clocks OK. He found B and C relays permanently operated in clock relay set as B22 and 23 H/R, thus preventing the release of C at next pulse. All clocks pulsing after cleaning contacts.

22 April :

Whitecroft in 2009 (picture from Google Earth)

Whitecroft in 2009 before the station was built (picture from Google Earth)

Before the clearances

Before the clearance work

After the clearances

After the clearance work

Paul and I went to meet Jason at Whitecroft to view progress on the preliminary work necessary before a cable trench can be dug. The work is well under way. We agreed a new DP3Y position and to a short part of the cable run being back on the fence as the gas main is obstructing a trench. We have asked for a loop to be left where the cable exits the ground so that a DP could be inserted later if the SPT layout at Whitecroft develops. Scott was obviously enjoying his big big machine.

Surplus power cable

We returned to Norchard to sort out the cables at the back of the caboosh. They were heavy. We moved the power cables out and now have a much tidier rear of the caboosh. We found a couple of significant lengths of telecoms cable.

After a Kaplan's lunch with Jean we sorted out surplus gear to take to Avoncroft tomorrow for the THG swapmeet. We are on our way to a tidier caboosh.

23 April : Paul had a table at the THG Avoncroft swapmeet. He made 98 for the fund. I thought it to be one of the best swapmeets.

26 April : Rick and Martin finally sorted out the line problem to signal post 2. Ray and Charles got the clocks all back to the correct time. I went to Whitecroft to measure the length of cable needed if we are asked to put a phone at the ground frame. 200m and 12m of trenching. Not too bad. I also sorted through the electrical stores with Alex, reducing the amount we hold to about half. I now need a trip to the dump. It's Thomas this weekend so Ray cleared up our room so that it is fit to be the ticket office.

May 2017

Whitecroft replacement cable

1 May : I went to see how the cable work at Whitecroft was getting on. The replacement cable is now in the trench apart from a 25m section near the gas main that needs us to tie to the fence line. A bit of gardening needed first to expose the fence. Scott has put the excess cable in their portacabin on site. He will also erect the wooden post needed for DP3Y to mount on. As there is no slack on the Norchard side of the DP, the cabinet needs to be remounted exactly where it is at present. Thanks Jason, Scott etc for your efforts in getting such a lot of work done to enable the trenching to take place.

3 May : Peter stayed at Norchard sorting the caboosh now that there is some space in there. Martin, Charles, Ray and I went to DP3Y at Whitecroft to find that the wooden post put in by Scott was exceedingly wobbly. We worked hard finding rocks to put around the post to stiffen it up. In the end we succeeded and the post may lean back a little but at least it seems to be stable. We got the cabinet off the old cables so that it could be drilled for fixing to the post. We also had to identify which cable emerging from the ground was the good one from Norchard. We got the final cables tied to the post and the cabinet screwed to the post. The very heavy new cable to DP3Z is threaded through the base of the cabinet but the Norchard side cable has to wait until we can cut the connections that still exist. A very heavy day's work for us and I am glad to see the cabinet on the post and ready for terminating.

6 May : Today was new volunteer's day. We had a presentation of the work covered by all the groups followed by a run up and down the line, a tour of the workshop and a buffet lunch. Following the presentation, Roger Hardiman from Coleford joined me to have a chat about telecoms and datacoms. He works in datacoms with a particular interest in CCTV and even has an interest in the Asterisk system. We visited Parkend and Lydney Junction exchanges and then had a look around the equipment at Norchard. Peter joined us for the Norchard tour as he was in for the day. Welcome aboard Roger, we hope to see you join us.

Whitecroft replacement cable ready to terminate

10 May : Martin, Ray, Charles and I went to Whitecroft and very quickly had the new cable ends stripped and ready for terminating next week. We were pleased at how quickly we managed to get the very thick sheath off the cable core. That confirmed the cable as having 14 pairs and using a railway colour code. After lunch we were joined by Peter when we had reports of both the BT line and the ground frame internal phone being faulty at Parkend. Ray and Martin found the ground frame wall phone damaged with evidence of an attempt to prise the case from the body. The phone has been put into a working condition but is rather wobbly on its fixings. The damage has been reported up the line. Peter and I found that there is a fault somewhere on our wiring from the BT master socket. There are six points radiating from the master socket to locations spread around the site. We need considerable time, and access to the ticket office when it is available, before we can continue faulting our wiring. It would be sensible for us to reduce this wiring to a minimum whilst we are working on the circuit. Charles went back to his bell circuits in the museum.

13 May : It was military day. 5541 was joined by Rennes and Sapper hauling passenger and freight trains around the railway. Rennes was officially welcomed back into service. Our day was largely spent familiarising Sam Bolinbroke and Roger Hardiman with the data equipment we use at the DFR and in having a meeting with Paul, Peter and myself to discuss forming a data group. Between them Sam and Roger seem to cover many of the aspects of data work that we need eg data, IT, CCTV. They agreed that they would accept data faults and that James etc should report them directly to Sam. Design work would be undertaken between them with help from Paul when required. We had some discussion on the projects that need to be pursued. Roger wishes to be involved in telecoms work when possible and Sam showed an interest in the strowger network. I managed to fit in a quick visit to Whitecroft to check on the colours in the new cable so that I can produce proposed DP cards for Tuesday.

16 April : Rick and I went in to changeover to the new cable at Whitecroft. As soon as we got there and took the cover off the first cabinet, it rained. It rained for the rest of the day so the job was abandoned until I return from holiday.

17 May : Another miserable wet day. I came in only briefly. Rick spent the day working on the 4000 type switch. After much swearing he managed to move the operating pin for the NP springs on to level 8 and adjusted it to operate the springs. Then he realised the vertical pawl hinge was seized and would not pivot on it's pin. Lubricated but not much movement - left until next week. Martin spent the day on the computer correlating our small stores - screws etc. Charles spent more time in the museum working on the token machine. Ray and Peter worked on the BT faulty extension wiring at Parkend. Eventually a fault was found in the cord connecting the filter to the LJU. A new cord was fitted and tested ok.

20 May : Peter, Sam, Roger, Rick and I met in the cafe. The main topic was the need to establish a data link from the shop to the cafe area inorder to sort out our Wi-Fi problem. We did a quick survey covering accommodation for a data rack and the route for CAT5 cables across the tracks. We will need to return using the tone set to find the old route used by the BT cable to the old kitchen car. The obvious place for a new data rack is in the shop store room. Sam has his proposal more or less ready to submit. Rick had a look at the 4000 type selector mechanism that has been soaking in oil. It's not free enough to put into service and we may need to use another mechanism or vertical magnet assembly. We have the spares.

Data racks arrive

24 May : The new Data Group made their presence felt today. Roger had organised the delivery of a couple of data racks recovered from one of his jobs. They arrived at Norchard. Luckily Sam had joined us for the day. We got one rack into the shop store room where we hope it can be installed permanently. The other was taken by train to Parkend where it now resides in the goods shed awaiting a passageway to be cleared into the exchange. Rick got on with reorganising DP3Z at Whitecroft. We are taking the oppotunity to really tidy up this DP before we change over to the new cable. After lunch, Sam, Peter, Ray and I went looking for the pipe that used to feed the BT line to the old kitchen coach from the shop to the platform. We found both ends but there is a join somewhere, now buried deep under the platform and we are not sure where. This involved lifting several large concrete tiles and having a dig around, thank goodness for young Sam. We have concluded that the pipe is a write off and that data between the shop and the platform will have to be carried via the telecoms cables using data extenders. Afternoon tea was well deserved.

27 May : Roger, Peter and myself went on the train to Parkend to get the data cabinet into the exchange and the battery racking back to Norchard on the train. The racking is now installed in the caboosh and already has some strowger and krone gear stowed on it. This gear was brought in by Roger as recovered from one of his jobs. Much of the day was spent on going over proposals for the first two data jobs we intend to do. We went to lunch with Jean at Kaplan's.

The first job is to provide Asterisk extensions at Parkend as a test bed for a VOIP network. We will attempt to do this job without involving the railway in any expenditure. We will need to run a group of external CAT5 cables from the router in the ticket office to the new rack in the exchange. There we intend to fit a switch and a group of ATAs. Analogue key phones will be connected to the ATAs via our cable network. I would expect key phones to work all round Parkend site but it will be interesting to see if we can extend an ATA to a key phone at Whitecroft, about 100 loop ohms away. All of these keyphones will be numbered in the 4** range. They will provide a useful alternative should the cable to Norchard be cut again.

The second job will be to reorganise the two routers that we have on the Norchard site. We intend the cafe router to provide the Wi-Fi facilities all over the site. The office router will handle all the office and shop data requirements. This will mean that Wi-Fi and data will be entirely separated. The main job will be to provide two data circuits between the shop and the cafe using our cable network and data extenders. We hope to place a further data cabinet in the station master's office as this seems to be designated the home for the Officer on Duty. In this case other data circuits may be needed from the station master's office and the hub will then be ready to grow. This proposal involves obtaining permission for the accommodation requirements plus a budget to purchase the gear.

31 May : Rick and Martin continued to organise DP3Z, Charles and Ray went gardening to find the fence line where we need to attach our new cable and Peter and I sorted some space in the caboosh and cleaned up over thirty Krone blocks we have been given.

June 2017

3 June : Paul, Peter, Rick and I went to Parkend. Rick and I looked at the ground frame phone and found the switch hooks fouling the case again. What do the signaller's get up to? Rick readjusted the hooks and when tested we could not make the phone fail. We checked the exchange equipment as we had a noisy line on one connection. We didn't find the fault but we did find a selector that had failed to release and had burnt out its release fuse. A problem for Ray on Wednesday. Paul and Peter stripped out items from the new data rack for transport back to Norchard. After lunch Paul, Peter and I had a walk round Norchard with Jason to hear what the likely plans are for the site. We will need some duct provided when Jason does trenching.

7 June : Rick and Martin worked on the DP3A and 3S jobs, getting the DPs ready for wiring. Ray cleared a fault on a jammed selector at Parkend. Charles spent most of his day talking to members of the public, particularly about his token machines in the museum. Peter worked on tidying stores. I had a very unproductive day, though I did find a 4000 type selector at Parkend without a seized vertical pawl for use as a donor in making up a spare demo selector for the museum.

10 June : Sam and Peter established that the pair providing the data link to the Junction signal box was "dis". It took most of the day to track it down to the krone block in DP3E. We need to get rid of that block and sort out the DP, but its a biggish difficult job on a very vital DP. It will have to be a crimp job as there is not enough room in the cabinet for connectors of any sort. Paul and I also managed to connect the green museum phone to the new demo first selector. I have started to build a further demo selector to act as a spare.

13 June : Rick and I met to get on with the job of changing over to the new cable at Whitecroft. It was more awkward than we first believed as the actual wire colours were nothing like the recorded colours, but with a lot of tapping out and getting pairs straight and labelled we then started to make progress on the changeover. DP3Z at Whitecroft now looks pretty tidy and just needs a bit of a clear up and some new labelling, DP3Y also needs tidying. Testing was pretty good, just one outgoing junction from Parkend to Norchard was faulty and we think that may be a fault unrelated to our job. Lots of midges to annoy us during the work.

14 June : After the usual tea and cakes- Ray and Charles went off to Parkend to check the clocks and leave a message for Stuart Berne. Rick and Martin went looking for a fault on I/C junction 3 from Parkend to Norchard (busied out yesterday). Testing int dis b wire toward PKD. At DP3A the d side jumper was testing HR b wire ( 34 volts) but the e side jumper tested 50 volts. The jumper was refitted and then tested OK. Peter was busy with some data work.
After lunch Ray and Charles returned to PKD to test the junctions to NCD. Both working OK. Ray went home and Charles returned to NCD to help Peter to measure and calculate runs for data cables. Rick and Martin went to Norchard signal box to check on a fault report. 525 had been reported faint I/c calls. The receiver inset was changed and tested OK with the signalman. Unfortunately signal post 2 telephone was permanently calling. The weatherproof phone was checked OK. The echo tester was used and showed a s/c 350 meters to the fault from the signal box. A walk along the line to check the cable was not helpful. As it was late afternoon it was decided to leave it for another day.

17 June : Peter is working on the 10+50 museum switchboard while Paul and I decided to tap out the 20 pair to DP3A and get it terminated. Not easy, the connectors we have are rubbish and probably should be changed out. We could not get reliable connections and then pair 3 was finally declared dead and gone as there was no white wire terminated on the frame. The tone machine indicates that it could be cut off in the butt, so we have a nineteen pair relief cable. I went off to Parkend to an S and T management meeting only to find no one there. I assume it was cancelled without us being told. We ended a not too brilliant day unable to repair a demo selector that was buzzing away on the 11th step. I think it's the wiper but we will wait for Wednesday to find out.

21 June : Rick and Martin checked on the PG SPT line at signal 13. Eventually they cleared the problem by changing the pair in use but did not really find the fault. Peter and Sam provided a new line to the undercroft for Permanent Way. Charles worked on his museum display. Ray went to the Junction to locate and bring back various items. I installed the new demo selector. It worked well first time, so now we have a spare. I also jumpered up 593 as the new line for Permanent Way.

28 June : The usual morning tea and cakes but with Sam and Alison. Milk by Charles cakes by Rick. Heavy rain is going to make working difficult today. Ray and Sam went off to Parkend to return a spare ringing machine (to be tested on a non running day). While there, they were told that calls to Norchard were not working. Test calls passed all ok. Possibly congestion!!! Caused by Roger Phelps testing the token from the ground frame and station staff trying to contact Norchard? Charles and Alison went off to meet their friends. Martin and Rick checked the exchange and found a PG alarm which turned out to be the extn on the high level platform. The handset was replaced - after a walk in the rain. They then worked on the 4000 type demo switch - wipers changed and adjusted - now available as a spare for the museum phones.
After lunch Ray and Sam continued work on the PA upgrade. Martin continued upgrading the computer records. Rick went home early to take his car in for service and mot.

July 2017

5 July : Only Rick, Martin and I were in today. Rick and Martin spent the day getting the new cabinet for DP3S ready for the changeover. We need to recover a length of ten pair from Whitecroft before we can make a start on the job. I spent the morning cleaning up a load of scrap copper by cutting off all the brass contaminants. After lunch I went to Bendall's with a modest load of scrap copper and cable. 51 for the fund.

8 July : Only Sam and Peter in today. They had a fault on the data link to the Junction to chase. In the end they swapped extenders with our office link and everything worked. The problem extenders also worked when tested back to back, so we feel we may have a problem with the sockets or similar. They also finished off fixing the cabling to the Undercroft phone.
In the evening the group went on the fish and chip special, a good trip.

12 July : Rick and Martin spent the day recovering a couple of lengths of the spare ten pair cable at Whitecroft. Unfortunately somehow this resulted in a dis wire on the Parkend alarm pair. Cutting the crimp off and making a new connection sorted that out. I am now having to think that crimps have joined krones in the unreliablity stakes. I hope we do not need to go back to the S&T links to make reliable connections. Peter and I chased an alarm from the Junction. We found the battery down to 36 volts. This was puzzling as everything checked out. However we finally came across a short circuit contact in the clock pulse circuitry which resulted in there being three relays permanently operated and a lamp lit rather than just one relay. With the primitive trickle charge arrangements in place we think the drain was too much for the charge and the battery slowly discharged down to 36 volts. We put the portable charger onto the battery and got it back up to normal by the end of the afternoon. We also chased a box to box circuit fault, eventually finding that the microphone battery at the Town station had failed and ruined the battery holder. Replacing the lot got the circuit up and running again, hopefully for the next ten years.

15 July : Roger, Peter and I spent the morning discussing the proposed Asterisk installations at Parkend and following this up with a site survey visit. Roger will obtain an ATA on a buy or return basis for us to try out at Parkend to see if it could be suitable as a standard item.

19 July : Rick and Martin put in the cable between the old and new DP3S, Peter and Charles sorted out some good phones for the stores and did some jumpering for the museum 10+50, and I went to check the battery at the Junction. It was fine, however the PA was not working due to a broken microphone lead. I decide it was not repairable so I ordered some replacements from the internet, 2.30 each, carriage paid, amazing.

22 July : Peter and Sam got on with providing three new krone strips mounted in the back of the 10+50 PMBX. This was made difficult by the cramped, dark working conditions at the back of the PMBX, cheap cable ties, and having to borrow tools because Peter had forgotten his! While they got on with that, Paul made some modifications to the asterisk server to prepare for testing at Parkend.
Roger is one of two new data technicians who have joined our group. He is familiar with our "Asterisk Exchange" which, with his expertise, we now feel can be brought into service use for the railway. This is an exchange within a computer. We run this equipment already and it is interconnected with our internal heritage system. It provides lines in the 4** range. It does this via data links and these links can go anywhere the internet is available. At present, members of the Telecoms Group have Asterisk lines at home, connected to the internal system via the internet. We mainly use these lines for checking the alarm system provided on the strowger network, but we can ring in to any where on the internal network. We therefore know that we can connect data keyphones to our internal network via data links.
In the autumn we intend to use Parkend as a test area for the provision of such lines from the Asterisk. All dial phones at Parkend will have a keyphone fitted alongside, and these will be connected back to Norchard via the internet. The dial phones and the keyphones will have identical facilities for making calls on the internal network and the users will have a choice as to which phone to use. One advantage from this work will be that Parkend will not be cut off entirely from the rest of the railway should the telecoms cable be damaged again.
If this work proves to be satisfactory to the users, we would expect to gradually extend the data circuits and keyphones over the rest of the railway, probably over several years. We could recover dial phones when the keyphones become sufficiently reliable. This will give us a route towards being able to dispense with the Strowger exchanges before we reach the point where we will have no one qualified to maintain them.

26 July : Rick and Martin changed a sticky dial at the Junction. Charles worked on his museum token bells which now seem to work. Peter did a bit more on the 10+50 . I went to check Parkend exchange and tell Andy at Parkend about our plans for keyphones at Parkend.

29 July : Sam, Peter, Paul and I were in though Sam left early. Paul worked in the exchange running cable for his speaking clock project. Peter and I surveyed Parkend for telecoms work needed for the Parkend VOIP experiment. Quite a list of jobs to be done.
Build a new control circuit and recover the old ringer/power control circuitry.
Recover the common service feeds that would have fed a concentrator in the signal box.
Provide a new krone MDF adjacent to the data cabinet.
Provide a 40 pair tie cable to the old MDF.
Shift the 20 pair cable to the signal box to the new MDF, straighten the pairs while also changing the DP in the signal box to a box conn 301A.
Shift cable from main building to the new DP while tidying the leads around the building.
Shift cable from goods office to the new DP.
Change Norchard admin office OP box 254A to a box conn 201D .
Mount two box conn 254A under new MDF in readiness for eventual transfer of 24 pair 0.9mm cable to DP1 to new MDF. These will be a joint between the 0.9mm cable and 0.5mm tails to the MDF.

Peter is now away for a couple of months and we can tackle these jobs in his absence.

August 2017

Bernard's gate to DP3S

2 August: Bernard Baldwin has built us a gate to DP3S. This will be most useful as we will no doubt need frequent access to this DP. Thanks Bernard.

Rick and Martin had a go at a weird fault apparently on a first selector which caused the call to be continually held, until of course you tried to find the holding condition. It looks as though there may be an HR contact somewhere. It didn't get found so it will be there for our next visit. Ray and I managed to get a 6*2 contiboard up to Parkend in the rain where it will be used as a backboard for the proposed new krone MDF. Charles had chats with the public in the museum. Not a lot accomplished today, but then the weather managed to depress us all.

5 August : Roger and I spent the morning checking the three internet connections we have at the railway, mainly to see what speeds we are getting. Parkend was the winner with a 74 Mb download speed and an 18 Mb upload speed. Can't do better as the cabinet is right there at the station. Norchard office was next with 32 Mb download and 5.8 Mb upload, and finally the cafe provided 28 Mb download and 3.9 Mb upload. We are looking good for all the projects we have in mind.

The data cabinet and new krone MDF at Parkend
9 August: We started the day with a teach in about the work required at Parkend for our experimental Asterisk extension project. Martin, Charles and I had an early lunch after getting our tools ready for work at Parkend. We then went to Parkend to install a backboard for the new MDF and to get the new krone base fitted. We got the tie cables to the old MDF and the earth cables to the consumer unit run in ready for termination. During the visit I also recovered the common service cable that was going to serve the signal box concentrator before we decided to fit the new exchange/concentrator there rather than at Whitecroft. A good day's work. Now we have to work out the best way to terminate cables on the new MDF. E sides from the cabinet at the top, then D side cables out to the network and the temporary tie cable to the old MDF at the bottom I suggest. How about that, Peter? Then we need a cable between the MDF and the cabinet and we have no idea where or how we terminate in the cabinet. Any ideas, Roger? At least the day finished with the sun shining.

12 August : I went up to Parkend for a quick visit to check for Peter that the wiring shown on the DP2 card is correct. It was. My guess is that Martin and I in 2006 terminated the DP cable at random one end and then tapped out to get it right at the other.

16 August :

new MDFOld MDF

Martin, Ray and I went to Parkend. We finished tieing in the old to new MDF cable and made a start in fanning out the wires and terminating. I am fighting the ordinary strip connection on the old MDF and Martin is having a go at the krone connections. Ray stayed back at base for the afternoon and got the office into shape for use as a booking office for the Thomas weekend.

Parkend new MDF

23 August: Rick and Martin did a little on DP3S and then concentrated on a fault in the Norchard exchange. They narrowed the problem down but at the end of the day had not been able to clear the problem. The linefinder and first selectors remain busied for now. At Parkend Ray and Charles ran in the cables between the cabinet and the new MDF. They stripped the cable ends at the MDF ready for terminating on krone strips. I completed the terminating of the tie cables at the old MDF. Very tedious to terminate wires on a 25 way wire wrapped block but it's done now. We have heard that the cables to the signal box will need to be cut off, pulled back and threaded through a new lead-in duct sometime in October. We will changeover the cable to the new MDF in Parkend exchange at the same time. We will reterminate the cables on a new krone Parkend North MDF in the signal box.

26 August : Paul and I spent the morning installing Paul's speaking clock relay set. It is working on 846 (TIM). Then we realised that only one call could access TIM at any one time, so we decided to build three more access circuits and jumper them to 847 to849. We could then PBX hunt over the group when necessary. This would be the only time we have used the PBX hunting facility. We picked up Sam on the way to Kaplan's for lunch. We then picked up relays and a line transformer so that Paul can enlarge his relay set.

After lunch we went to Parkend to let the data boys see how we are progressing with the asterisk rearrangements. They were happy with our progress. We discussed the cabinet layout and the means by which we will connect the ATAs to the line pairs. We concluded that the way Paul connects such circuits was the tidiest. We will terminate the tie cables on to an RJ45 patch panel in the cabinet, connecting 4 pairs to each jack. Paul will make up some cords that will connect the 4 way RJ45 jack to four separate one pair cords terminated with RJ11 plugs. Each plug can then plug into an ATA line output jack. Very simple, very tidy, minimum problem points, sounds good to me.

We then considered how to move the router from the office to the exchange. We got lost for some time trying to work out how the office had been wired. The real problem was that the BT extension wiring had been taken to a socket serving an internal line so that the BT line could get access back to the exchange in the same cabling. We came to the conclusion that we must replace all the office cabling and make a decent job of it before we can shift the router. Peter to note - you were right, the wiring is rubbish.

30 August : Rick spent further time on the tricky intermittent fault on rack A2. Eventually it vanished before being found. Annoying. Martin finished terminating the old to new MDF tie cable at Parkend. Charles terminated the 40 pairs of the data cabinet to MDF tie cable onto a patch panel in the cabinet. Useful practice as no doubt we have made a hash of the job as we have no experience of such work. Plenty of cable length when it gets reterminated!!! Rick brought the faulty first selector back to Parkend. It's fine but testing revealed a fault on a final selector. No battery to the A relay, a fault harking back to the time when we were going to run the exchange straight from the mains and start it with a 12 volt battery. I need to demodify the finals so that the battery supply is more straightforward. As we were leaving I was introduced to a possible new recruit, Tony Williamson. He has been a lineman among other jobs, so he could be useful if he decides to give us a go.

September 2017

2 September : Paul, Roger and I went to Parkend. We seemed to have some problems on the exchange when we put the call sender on, but all problems vanished into thin air. I mounted some components to build a ringer control circuit in place of the current over complicated relay set which I hope to recover. We traced wires to find that LR line relays and A relays on the incoming finals got their battery from a separate fuse. Part of the problems from the original design spec for the exchange, this will not be a current problem but it could be demodified in the future. For now it needs to be better explained on the wiring diagrams.
Paul and Roger inspected the work done so far on the analogue RJ45 jacks and MDF strips and pronounced themselves to be happy with the work. Very satisfying for our untrained telecoms volunteers. Paul and Roger tied the cabinet to new MDF cables into place, cut off the excess and made a start on terminating on the MDF, leaving some work for our telecoms guys to complete. They have decided that the analogue side of the equipment is to be mounted at the very top of the cabinet.
Lunch at Kaplan's with Jean as usual.

The and new krone MDF at Parkend
6 September:


Rick and Martin did a further survey of DP3S and then faulted a selector in Norchard exchange. It turned out to be a pair of selector U points shorted out. They changed the U points but have still to repair the wiring damage they caused during fault finding.

Charles and I went to Parkend. I labelled much of the exchange, a job long overdue. Charles terminated the data cabinet to MDF cables on the new MDF. We tapped them out and also the tie cables to the old MDF to find no problems. We also mounted an extra 6' fluorescent fitting in the exchange in the hope that we will get more light to the new MDF. Alex has said he will get it wired up when he provides the cabinet power.

9 September : Paul and Roger ran in some cable at Norchard for Paul's expanded speaking clock relay set. His set was checked as working on its original single exchange line. It will now cater for four lines in a PBX hunting group, but awaits cable terminating and jumpering next week. We had a very short working day and left before lunch.

13 September : A quiet day with just three of us in. Rick spent the day on faults in Norchard exchange. Charles and I took scrap cable and copper to Bendall's and then went to Parkend to partially label the new MDF and install the jumper rings. The MDF is beginning to look busy.

16 September : Paul, Roger and I terminated the wiring for Paul's speaking clock relay set, got it jumpered up and tested. Some minor problems were quickly sorted and the set tested out correctly. Then we found an odd design error. If more than one person dials TIM simultaneously, then they can speak to each other over the clock. Probably this will never be a problem in the museum where everybody is next to each other in any case, but it might be difficult should TIM become used by the rest of the staff. One limitation is that Parkend staff cannot dial 8, so cannot access the clock.

We also surveyed the route for the CAT5 cables to Parkend office and found that we are short of external grade CAT5 and can only provide two cables at the moment. The data boys prefer buying a small switch for the office to spending money on another reel of external CAT5. We will see, we cannot find a non running day on which to do the job until November in any case, and perhaps by then we will have more access to cash.

20 September : The faulty group selector from last week was repaired and put back in service. Rick, Ray and Charles installed a new 6 wire to the Norchard crossing hut phone in preparation for the cable changeover at DP3S. We also cut and stripped the new 10 pair cable and left it covered and sealed ready to go in DP3SA. Charles and I went to Parkend to look at the token battery but there was no locker key available. We did not get access in the end but Dave Collins let me have a key for a copy (or two) to be made. Peter is back from Pakistan and came briefly to Parkend to see progress on the Asterisk project. He is still not over the flight and felt awful so went off home early. Charles and I watched the signals operation at Parkend ground frame when a two train operation is underway. Ray and I left at lunchtime. Charles delivered the token cabinet key to my door at the end of the afternoon, but the key cutting shop was closed when I got there. Afternoon tea was abandoned as the cafe was so full.

Soldering up a set of token batteries
23 September:


Peter came in briefly. Paul and I spent the day on token batteries now that I have token cabinet keys available. We tested the Parkend 48 volt set to find them dropping below 30 volts when asked to ring the distant bell. We decided to change the set which cleared up all the problems. The set was only 18 months old, so we concluded that we are running more trains on so many more days that the battery life is being affected. We decided to check the three other token battery sets to find them all beginning to show signs of failure. We will need to change all the batteries on Wednesday. We made up two sets and have sufficient batteries available for the final set. It looks as though we should now change the token batteries annually rather than just checking them every other year as in the past. Paul and I met Jean for a Kaplan's lunch.

25 September : The phone rang. Jean said "It's Chris Bull for you". That meant only one thing, a cable had been cut. I called Rick and we walked along north from Upper Forge to DP3W where we found the cabinet completely missing and the cable ends stuck out of the ground. Nothing could be done there. We returned to Norchard to see what jointing options we had. We found a compression joint which we could try. We disconnected the incoming Parkend junctions and I went to Parkend to find the exchange ringer running continuously. In the end I simply switched the exchange off to stop it running to no purpose. We hope to start a repair tomorrow.

26 September :


Rejointing DP3Wfour red lights on STARS all at once

Rick and Martin had got the joint made when I joined them and we are ready to concrete in a post to mark the joint. Whilst coming back to Norchard for lunch I got a text saying that Chris Bull had cut the cable to the south at the Town crossing. We were joined by Charles and we had a quick lunch in the cafe. Then we went to see the damage. Much larger as it is a 50 pair armoured cable we are dealing with. It had been a drainpipe joint but it was entirely broken up. We will need an S and T cabinet inorder to make a repair. I left Rick and Martin to get the cable ends into plastic bags while I went to the Junction signal box to disconnect the exchange. This is turning into the worst week of the year.

27 September : The railway is running just a single train today with pilotman working! I went to Parkend to switch the exchange back on and get the junctions tested. All was well and the exchange is functioning satisfactorily again. We could not test the token as no signalmen were available. The Lydney Junction cable fault was not tackled as we needed to come up with a plan. We decided that an S and T cabinet was not the answer as there is insufficient space for one between the bank and the rail. Our lineside group did not want us to replace it as a drainpipe joint, seeing that as a potential for further damage next year. I managed a chat with our new General Manager. We decided to do a really quick and very temporary repair to simply put through the few main circuits we need. We will end up with the cable ends left in a stout plastic bag. We will then find, or even purchase, a replacement length of cable to go from cabinet to cabinet, bypassing the narrow danger spot, thus eliminating the problem. Further he will call a meeting of interested groups to see how we could avoid these problems in the future. Rick and Martin also arranged for postcrete to be delivered to the Parkend cable fault site along with a visibility post for erection tomorrow. Charles put together a further set of token batteries and fitted two of the sets in Norchard box. We still have Lydney Junction box in which to fit a new set.

28 September : Rick and Martin went up the line to concrete in the marker post for DP3W. Then we went back to the Lydney Junction cable problem. Rick and Martin concentrated on finding the important circuits from among the debris caused by the flail. I lost count of the running up and down to Norchard and Lydney Junction I did in trying to help. At one stage even Jean walked up to the fault with me when it was time for lunch. We ended the day with one junction outgoing from Lydney signal box, two incoming circuits and the token. At least we hope that the token will work as the galvos deflect. That should allow the weekend to go ahead as planned. Thanks Rick and Martin, a super effort.

October 2017

4 October : The day started with tea/coffee and cakes provided by Janet. While discussing the day's work the lads were visited by Roger Phelps who said he wanted to thank 'The Team' for their excellent work last week.
It seemed like an Awayday to Parkend for 'The Team' as they all had things to do that way. Charles went to check the token batteries - Ray and Peter went to check the exchange and survey the signal box cables - Rick and Martin went looking for the damaged drop wire. During lunch Sam arrived looking to help out. After lunch Ray went home as he did not feel well. Charles visited Norchard signal box for some more experience , Peter and Sam returned to Parkend and Rick and Martin returned to complete the repair to the drop wire serving 304 - the telephone in the box at the signal on the Parkend straight. They all returned for afternoon tea at Norchard where they met Steve Harris and talked about the future work required at Parkend signal box.

4 October : Peter went to look at the new MDF at Parkend and realised that there were reversals in the new to old MDF tie cable. Sam and Peter spent most of the day sorting this out by shifting the cable so it is now run inside the krone frame and completely reversal free.

6 October : Peter met Sam at Parkend in the afternoon and recovered the BC 254A in the signal box leaving the S/B end un-terminated with everything disconnected. They then proceeded to the exchange, and removed the cable from the old MDF and pulled it back stripping the sheath ready for terminating.

7 October : After the usual cup of tea and a catchup, Peter and Sam went to Parkend to terminate the cable that goes to the signal box. One end is now neatly terminated on the new Krone MDF, and the other is ready and waiting to move it to its new location. Paul and Roger busied themselves troubleshooting the UPS that powers the Asterisk. After taking it to pieces, and testing it thoroughly it looks like the battery is fine but there is a fault with the inverter. The battery charges ok, and holds a charge, but when the mains goes off the inverter doesn't start up to feed mains to the Asterisk. Fixing this was deemed to be more trouble than it's worth. The Asterisk was reconnected directly to the mains supply to restore service, but who knows how long that will last on Norchard power. Steve Harris was in the Parkend box when Peter arrived and asked if he could remove all the phones from the wall board, this was done along with this all the old wiring being removed. The wall board is now ours and is now in the locking room. It needs refurbishment before reuse.

8 October : Peter spent Sunday afternoon in the workshop making a simple jig to align the Parkend MDF jumper hooks so as they would be the same as the jumper rings in a customers premises.

9 October : Peter met Sam at the Postage Stamp cafe for tea at 10:00 o'clock and proceeded to Parkend armed with the hook aligning jig and a drill. The hooks were refitted and look the part. The jig is suitable for any frame 50 - 340 pairs. Having now got the frame almost complete for use they started to change over the existing S/B circuits checking the polarity as they went. Along with 3 extensions they also connected through 3 junction circuits using the correct jumper colours.

11 October : A rainy day. Roger Phelps popped in again to thank the group for the effort made in getting service back after the flail damage. He also obtained authority to purchase the required replacement cable when we were ready to install it. Peter, Ray, Martin, Charles and myself had a long chat about where we were at Parkend and then went to Parkend to look on site at the situation. Peter brought everyone up to date. The signal box is progressing and we can move in the exchange and concentrator rack any when we like. At least the rack can be mostly taken apart for transport so that is a job I can get on with. We agreed what phones we will provide in the box ie the heritage concentrator phone on the wall beside the concentrator head, a table 746 for 725 and two MF phones for BT and the Asterisk. These phones will simply sit on the top of the cupboard. During this we realised that there was a low voltage alarm on Parkend exchange. We investigated to find that the charge circuit had been switched to the "other" rectifier which of course was off. We cannot think how this occurred.

Parkend North waiting for transport
12 October : Now that Parkend signal box has been tidied up and is able to receive the exchange and concentrator, I have dismantled the rack in my garage so that it can be moved as soon as we can find some transport.

14 October :


The damage to Parkend DP1EThe damage to Parkend DP1E

Jean and I went to Kaplan's to meet Peter and Roger for lunch. After lunch Roger and I inspected the damage to DP1E at Parkend. Jason had backed the digger into it and had done an excellent job of writing off the cabinet and concrete supports. The token, the alarm pair and a junction in each direction are not working. We went on to Parkend to discuss a new cabinet with Steve Harris. If we get one it will be enormous. As we only have a short length of cable to go into the cabinet, I can see us laying on the ground to do the jointing. Separately I delivered the concentrator to the locking room. Peter, Roger and I had a chat about the upcoming requirements for the Parkend Asterisk job. I have a list of bits we need. I left Peter and Roger discussing the work and went on to Norchard to a meeting to discuss lineside clearance etc.

Paul Smith was chairman, others attending were Chris Bull, Jason Shirley, Alex Davies (permanent way), Adrian Copley (S and T) and myself. Much of the discussion was about clearing the lineside and getting staff trained to operate diggers, cranes etc. I was tasked to survey the line for telecoms equipment and 1) to record the position of our plant and cables onto a large scale map as best that I can and 2) to produce a photographic file, one page per item, with any necessary comments of all our above ground equipment. These pages will be integrated into a comprehensive file of all lineside equipment that will be the basis for a regular spraying treatment. These pages will simply be an update of our present DP safety records.

Jason will also lay out and dig in, where possible, the 400m of cable to be ordered for the Town Crossing renewal job. I feel though that we have service to Lydney Junction and that we should firstly concentrate on the Parkend cable fault.

It was agreed that all lineside work involving construction or recovery changes will be notified to Chris Bull. I suggest we send him an email if we propose to change our cabinets, joints etc.

Chris Bull wants all drain pipe and other joints made more obvious and safe by the provision of a pair of stout uprights of some description protecting our joint. We will need to produce a policy for marking our less obvious lineside equipment.

18 October :


Parkend DP1ETemporary

A bit drizzly today. Sam, Martin, Charles and I went to Parkend DP1E to see what could be done following the major damage of last Saturday. We got the cabling clear of the broken cabinet and had the cabinet moved to the railside. Then we started checking pairs, in E side order. We found all the wires in the end quite easily as most of them were still being fed through the holes in the fanning strip. We made the new connections for all the circuits except the S and T wires going away to the adjacent signal. They must wait for S and T to come and reconnect them. (Subsequently I heard that Peter Wood will be in on Sunday to make these temporary connections) Sam went off to test the token. Thank goodness it worked after we found a couple of dodgy connectors. We finished with the joint covered temporarily with some strong plastic bags, not red, as that apparently can upset enginemen. Sam got the train to stop on its way back to Norchard and we loaded the damaged cabinet onto it for transport back to base. After lunch, the lads got the cabinet off the platform and down to the workshop while I returned to Parkend to check the junctions. All working it seems, hope everything else is! We had a discussion with Bob Hawker about renovating the cabinet. He is considering the best way forward, hoping to come up with a new cabinet. We also had a discussion with Peter about the extent of data work that we will undertake along with providing the Asterisk phones at Parkend.

We hope to move the Parkend North rack from my home to Parkend signal box next Wednesday. Keith Wakely will assist by providing a trailer on which to move the gear. We meet at ten for tea and coffee and then move on quickly to my house.


Current Value of our Fund

38.93


Return to the Diary Menu


mainmenu

Page provided by John Bathgate

This page was last updated on
20th October 2017